OpenStudy (anonymous):

what is simplifey 10/20 what is simplifey 10/20 @Mathematics

OpenStudy (anonymous):

2

OpenStudy (anonymous):

thanks 

OpenStudy (karatechopper):

1/2

OpenStudy (anonymous):

ok what about 9/12

OpenStudy (anonymous):

if they are looking for a whole number divide 20 by 10 if they are looking for a fraction then karatechopper is correct in the 1/2

OpenStudy (karatechopper):

3/4

OpenStudy (anonymous):

you're just dividing the numbers. put the numbers into your calculator \[9\div12=\] 0.75

OpenStudy (karatechopper):

i thought he was simplfying

OpenStudy (anonymous):

i am

OpenStudy (anonymous):

always divide by the common denominator. 9 and 12 can both be divided by 3 so (9/3)/(12/3) = 3/4=9/12

OpenStudy (anonymous):

so simplifying 9/12 is indeed .75 or 3/4

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

iimka done thanks u guys

OpenStudy (anonymous):

i will be on to morrow

OpenStudy (anonymous):

np!

OpenStudy (anonymous):

C u tomorrow